OpenStudy (anonymous):

Part 1: Describe the graphs of the functions f(x) = 2^(x) – 3 and g(x) = –2(x) – 3. Part 2: Compare and contrast the domain and range of f(x) and g(x).

OpenStudy (anonymous):

\[f(x)=2^{x}-3 \]

OpenStudy (anonymous):

\[g(x)=2^{x}-3\]

OpenStudy (anonymous):

Will give medal!?!?!?

OpenStudy (anonymous):

@uri can you help me with this

hero (hero):

\(f(x)\) and \(g(x)\) do not appear to be the same.

OpenStudy (anonymous):

i mean g(x) in front of the 2 is - so -2

hero (hero):

\(f(x) = 2^x - 3\) While \(g(x) = -2(x) - 3\). That's what I see.

OpenStudy (anonymous):

Yes that is the problem.

hero (hero):

Okay, so if that's the case, one is exponential and the other is linear.

OpenStudy (anonymous):

Ok so part 1: f(x):linear g(x):exponetial ?

hero (hero):

Well, that would only be the initial part of the description. To describe them fully, you have to describe specific characteristics of the actual graph of both functions.

OpenStudy (anonymous):

i thought there is no exponetial? iSN'T THERE ONLY: Absolute value function, quadratic function, cubic fuction, radical function, rational function, and linear function?

hero (hero):

What do you mean there's no exponential function? How could there not be one?

OpenStudy (anonymous):

Lol nv my bad

hero (hero):

Some of the top math sites: Wolfram Alpha, Purple Math, and Wikipedia all give descriptions about them.